# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 4.3.2 RL for Formal Mathematical Reasoning
Formal mathematical reasoning refers to reasoning carried out in a formal language with precisely defined syntax and semantics, yielding proof objects that are mechanically checkable by a verifier. This paradigm is particularly suited for advanced tasks such as automated theorem proving (ATP) [316], where an agent, given a statement (theorem, lemma, or proposition), must construct a proof object that the verifier accepts, thereby ensuring machine-verifiable correctness. From a reinforcement learning perspective, formal theorem proving is commonly modeled as a Markov Decision Process (MDP): proof states transition via the application of tactics [^1] In Lean-style Interactive Theorem Provers (ITPs), a tactic is a command or small script that instructs the system to refine the current proof goal, with the resulting proof term checked by the ITP kernel for correctness., each of which is treated as a discrete action in RL-based proof search.  [317]. Under this formulation, formal theorem proving can be cast as a search problem over a vast, discrete, and parameterized action space.
Formal proofs are verified by proof assistants such as Lean, Isabelle, Coq, and HOL Light. These systems, often referred to as Interactive Theorem Provers (ITPs), deterministically accept or reject proof objects producing binary pass/fail signals as the primary reward for RL training, while some works also explore leveraging error messages as auxiliary signals [318, 319].
@@
The outcome-only paradigm was demonstrated at scale in 2024 with DeepSeek-Prover-v1.5 [316], which releases an end-to-end RL pipeline in Lean based solely on binary verifier feedback, resulting in significant improvements in proof success on benchmarks like miniF2F [320] and ProofNet [321]. The authors propose a variant of MCTS, i.e. RMaxTS, that incorporates intrinsic rewards for discovering novel tactic states to encourage diversity of proof exploration during inference-time search and mitigate the sparse-reward issue. Building on this direction, Leanabell-Prover [322] scales up DeepSeek-Prover-v1.5 by aggregating an expansive hybrid dataset of statement-proof pairs and informal reasoning sketches from multiple sources and pipelines such as Mathlib4 [323], LeanWorkbook [324], NuminaMath [325], STP [326], etc., covering well over 20 mathematical domains. This broad coverage mitigates the scarcity of aligned informal-to-formal (NL to Lean4) training examples, which are crucial for bridging natural-language reasoning and formal proof generation. At the same time, Kimina-Prover [327] Preview further emphasizes the critical challenge of aligning informal and formal reasoning. It implements a structured “formal reasoning pattern,” where natural-language reasoning and Lean 4 code snippets are interleaved within thinking blocks. To reinforce this alignment, the output is constrained—to include at least one tactic block and to reuse no less than 60% of the Lean 4 snippets in the final proof, ensuring close correspondence between internal reasoning and formal output. A recent work, Seed-Prover [328], integrates multiple techniques. It first adopts a lemma-centered proof paradigm, which enables systematic problem decomposition, cross-trajectory lemma reuse, and explicit progress tracking. It then enriches RL training with a diverse prompting strategy that randomly incorporates both informal and formal proofs, successful and failed lemmas, and Lean compiler feedback, thereby enhancing adaptability to varied inputs. At inference, it employs a conjecture–prover pipeline that interleaves proving conjectures into lemmas and generating new conjectures from the evolving lemma pool, substantially improving its capacity to tackle difficult problems. Complementarily, the accompanying Seed-Geometry system extends formal reasoning to geometry, providing state-of-the-art performance on Olympiad benchmarks. Together, these efforts demonstrate that sparse but explicit reward signals can yield nontrivial gains, particularly when paired with effective exploration strategies.
@@
To improve credit assignment and reduce wasted exploration, several works extend the outcome-only paradigm with denser, step-level signals. DeepSeek-Prover-v2 [329] designs a dual-model pipeline to unify both informal (natural-language) and formal (Lean4) mathematical reasoning to reinforce the proving reasoning ability. It introduces subgoal decomposition, where a prover model solves recursively decomposed subgoals and receives binary Lean feedback at the subgoal level, effectively providing denser supervision and improving both accuracy and interpretability. Following this dual-role collaborative mindset, ProofNet++[318] implements a neuro-symbolic RL framework featuring a Symbolic Reasoning Interface, which maps LLM-generated reasoning into formal proof trees, and a Formal Verification Engine, which verifies these proofs with Lean or HOL Light and routes error feedback back to the LLM for self-correction. Leanabell-Prover-v2 [319] integrates verifier messages into reinforcement updates within a long CoT framework, enabling explicit verifier-aware self-monitoring that stabilizes tactic generation and reduces repeated failure patterns.
@@
Although both outcome-only and process-aware reward paradigms have demonstrated encouraging advances, the scarcity of high-quality theorem-proving data further amplifies the challenges of reinforcement learning under sparse rewards as well as the design of step-level preference signals [330, 331, 326]. To mitigate these limitations, a prominent line of work adopts expert iteration (ExIt) [332], a framework that combines search with policy learning. This paradigm provides an alternative to outcome-only or process-aware RL, alleviating data scarcity by producing high-quality supervised trajectories. Instead of directly optimizing against sparse verifier signals, ExIt performs search-guided data augmentation: valid proof trajectories discovered by search and checked by a verifier are reused as expert demonstrations in an imitation-learning loop. It usually employs a two-role system: the expert collects valid and progressive trajectories via MCTS under outcome-only verifier feedback, while the apprentice trains a policy on these process-level trajectories and then shares the improved policy back with the expert, thereby bootstrapping subsequent rounds of search and accelerating convergence. Polu and Sutskever [333] introduces ExIt into formal theorem proving, demonstrating that search-generated expert data can bootstrap models toward tackling complex multi-step proving challenges. Later works adapt this design to Lean and other ITPs.
@@
When applying to formal theorem proving, naive tree search methods often face severe search space explosion when navigating the vast parameterized tactic space. To mitigate this, InternLM2.5-StepProver [334] introduces a preference-based critic model, trained with RLHF-style optimization, to guide expert search, effectively providing a curriculum that directs exploration toward problems of suitable difficulty. Lean-STaR [335] further enhances ExIt by integrating Self-Taught Reasoner (STaR) [336]. It first trains a thought-augmented tactic predictor on synthesized (proof state, generated thought, ground-truth tactic) triples. Then, in the expert-iteration loop, the model produces trajectories that interleave thoughts with tactics; trajectories with tactics successfully validated by Lean are retained and reused for imitation learning. Empirically, the inclusion of thoughts increases the diversity of exploration in the sample-based proof search.
@@
A recent work, STP [326], points out that solely relying on expert iteration will quickly plateau due to the sparse positive rewards. To address this, it extends the conjecturer–prover self-play idea from Minimo [337] to practical formal languages (Lean/Isabelle) with an open-ended action space and starts from a pretrained model. STP instantiates a dual-role loop in which a conjecturer proposes statements that are barely provable by the current prover, and a prover is trained with standard expert iteration; this generates an adaptive curriculum and alleviates sparse training signals. Empirically, STP reports large gains on LeanWorkbook [324] and reports competitive results among whole-proof generation methods on miniF2F [320] and ProofNet [321].
@@
ARTIST [91] Outcome -
[…]
Prover-Verifier Games (legibility) [310] Outcome -
[…]
START [312] Process -
LADDER [313] Process -
SWiRL [314] Process -
[…]
@@
[…]
ProofNet++ [318] Process -
[…]
@@ 4.4 GUI Agent
GUI agents have progressed through distinct training paradigms. Early systems used pre-trained vision–language models (VLMs) in a pure zero-shot fashion, mapping screenshots and prompts directly to single-step actions. Later, SFT on static (screen, action) trajectories improved grounding and reasoning, but was limited by scarce human operation traces. Reinforcement fine-tuning (RFT) reframes GUI interaction as sequential decision-making, allowing agents to learn via trial-and-error with sparse or shaped rewards, and has advanced from simple single-task settings to complex, real-world, long-horizon scenarios. Table 7 presents the majority of works studied in this section.
@@ 4.4.1 RL-free Methods
Early GUI agents directly leveraged pre-trained Vision–Language Models (VLMs) in a pure zero-shot manner, mapping screenshots and prompts to single-step actions without any task-specific fine-tuning. Representative systems include MM-Navigator [338], SeeAct [339], and TRISHUL [340], which differ in interface domains or parsing strategies but share the same reliance on off-the-shelf VLMs. While showcasing the generality of foundation models, these approaches suffer from limited grounding accuracy and reliability, restricting their applicability to complex tasks [341, 342].
@@
The SFT paradigm adapts pre-trained vision–language models to GUI tasks by minimizing cross-entropy loss on offline (screen, action) pairs, without online interaction. InfiGUIAgent [343] employs a two-stage pipeline that first improves grounding and then incorporates hierarchical and reflective reasoning. UI-AGILE [344] enhances supervised fine-tuning by incorporating continuous rewards, simplified reasoning, and cropping-based resampling, while further proposing a decomposed grounding mechanism for handling high-resolution displays. TongUI [345] instead emphasizes data scale, constructing the 143K-trajectory GUI-Net from multimodal web tutorials to enhance generalization. While differing in focus, these approaches all face the limitation of scarce human operation traces.
@@ 4.4.2 RL in Static GUI Environments
In static settings, reinforcement learning is applied on pre-collected datasets with deterministic execution traces, using rule-based criteria for outcome evaluation in the absence of live environment interactions. GUI-R1 [346] adopts an R1-style reinforcement fine-tuning pipeline over a unified action schema, using simple format and correctness rewards to improve step-level action prediction with modest data. UI-R1 [347] applies group-relative policy optimization to stabilize policy updates and improve exact parameter matching through a compact action interface and reward shaping for action-type and argument accuracy. InFiGUI-R1 [348] introduces a two-stage training paradigm that first distills spatial reasoning to enhance grounding, followed by reinforcement learning with sub-goal supervision and recovery mechanisms to improve long-horizon reasoning. AgentCPM-GUI [349] combines grounding-aware pre-training, supervised imitation, and GRPO-based reinforcement fine-tuning with a concise JSON action space, reducing decoding overhead while improving robustness on long-horizon sequences. UI-Venus [350] is a multimodal screenshot-based UI agent fine-tuned via RFT with custom reward functions and a self-evolving trajectory framework, achieving new state-of-the-art in both UI grounding and navigation.
@@ 4.4.3 RL in Interactive GUI Environments
In interactive settings, reinforcement learning agents are optimized through online rollouts in dynamic environments, requiring robustness to stochastic transitions and long-horizon dependencies. WebAgent-R1 [95] conducts end-to-end multi-turn reinforcement learning with asynchronous trajectory generation and group-wise advantages, improving success on diverse web tasks. Vattikonda et al. [351] study reinforcement learning for web agents under realistic page dynamics and large action spaces, highlighting challenges in credit assignment and safe exploration. UI-TARS [352] integrates pre-training for GUI understanding with reinforcement learning for native desktop control, coupling milestone tracking and reflection to enhance long-horizon execution. DiGiRL [353] introduces an offline-to-online reinforcement learning pipeline on real Android devices, combining advantage-weighted updates, doubly robust advantage estimation, and instruction-level curricula to cope with non-stationarity. ZeroGUI [354] automates task generation and reward estimation with a vision-language evaluator, then applies two-stage online reinforcement learning (training on generated tasks followed by test-time adaptation) to reduce human supervision. MobileGUI-RL [355] scales training on Android virtual devices with trajectory-aware GRPO, a decaying efficiency reward, and curriculum filtering, improving execution efficiency and generalization while keeping the system practical for large rollout volumes. ComputerRL [356] introduces an API-GUI hybrid interaction paradigm paired with a massively parallel, fully asynchronous RL infrastructure and the novel Entropulse training strategy—alternating RL with supervised fine-tuning—to empower GUI-based agents to operate efficiently and scalably in desktop environments.
@@
[…]
TRISHUL [340] Vanilla VLM - -
[…]
@@
[…]
WebAgent-R1 [95] RL Interactive -
Vattikonda et al. [351] RL Interactive -
[…]
MobileGUI-RL [355] RL Interactive -
ComputerRL [356] RL Interactive -
4.5 RL in Vision Agents
RL has been applied to a wide range of vision tasks (including, but not limited to, image / video / 3D perception and generation). Since the number of related papers is substantial, this section does not aim to provide an exhaustive overview; for a more comprehensive survey on RL for various vision tasks, we refer readers to two dedicated surveys in vision [357, 197].

### v4

@@ 4.3.2 RL for Formal Mathematical Reasoning
Formal mathematical reasoning refers to reasoning carried out in a formal language with precisely defined syntax and semantics, yielding proof objects that are mechanically checkable by a verifier. This paradigm is particularly suited for advanced tasks such as automated theorem proving (ATP) (Xin et al., 2025), where an agent, given a statement (theorem, lemma, or proposition), must construct a proof object that the verifier accepts, thereby ensuring machine-verifiable correctness. From a reinforcement learning perspective, formal theorem proving is commonly modeled as a Markov Decision Process (MDP): proof states transition via the application of tactics [^2] In Lean-style Interactive Theorem Provers (ITPs), a tactic is a command or small script that instructs the system to refine the current proof goal, with the resulting proof term checked by the ITP kernel for correctness., each of which is treated as a discrete action in RL-based proof search (Wu et al., 2021). Under this formulation, formal theorem proving can be cast as a search problem over a vast, discrete, and parameterized action space.
Formal proofs are verified by proof assistants such as Lean, Isabelle, Coq, and HOL Light. These systems, often referred to as Interactive Theorem Provers (ITPs), deterministically accept or reject proof objects, producing binary pass/fail signals as the primary reward for RL training, while some works also explore leveraging error messages as auxiliary signals (Ambati, 2025; Ji et al., 2025).
@@
The outcome-only paradigm was demonstrated at scale in 2024 with DeepSeek-Prover-v1.5 (Xin et al., 2025), which releases an end-to-end RL pipeline in Lean based solely on binary verifier feedback, resulting in significant improvements in proof success on benchmarks like miniF2F (Zheng et al., 2022) and ProofNet (Azerbayev et al., 2023). The authors propose a variant of MCTS, i.e. RMaxTS, that incorporates intrinsic rewards for discovering novel tactic states to encourage diversity of proof exploration during inference-time search and mitigate the sparse-reward issue. Building on this direction, Leanabell-Prover (Zhang et al., 2025k) scales up DeepSeek-Prover-v1.5 by aggregating an expansive hybrid dataset of statement-proof pairs and informal reasoning sketches from multiple sources and pipelines such as Mathlib4 (The mathlib Community, 2020–2025), LeanWorkbook (Ying et al., 2025), NuminaMath (Li et al., 2024a), STP (Dong & Ma, 2025), etc., covering well over 20 mathematical domains. This broad coverage mitigates the scarcity of aligned informal-to-formal (NL to Lean4) training examples, which are crucial for bridging natural-language reasoning and formal proof generation. At the same time, Kimina-Prover (Wang et al., 2025a) Preview further emphasizes the critical challenge of aligning informal and formal reasoning. It implements a structured “formal reasoning pattern,” where natural-language reasoning and Lean 4 code snippets are interleaved within thinking blocks. To reinforce this alignment, the output is constrained—to include at least one tactic block and to reuse no less than 60% of the Lean 4 snippets in the final proof, ensuring close correspondence between internal reasoning and formal output. A recent work, Seed-Prover (Chen et al., 2025c), integrates multiple techniques. It first adopts a lemma-centered proof paradigm, which enables systematic problem decomposition, cross-trajectory lemma reuse, and explicit progress tracking. It then enriches RL training with a diverse prompting strategy that randomly incorporates both informal and formal proofs, successful and failed lemmas, and Lean compiler feedback, thereby enhancing adaptability to varied inputs. At inference, it employs a conjecture–prover pipeline that interleaves proving conjectures into lemmas and generating new conjectures from the evolving lemma pool, substantially improving its capacity to tackle difficult problems. Complementarily, the accompanying Seed-Geometry system extends formal reasoning to geometry, providing state-of-the-art performance on Olympiad benchmarks. Together, these efforts demonstrate that sparse but explicit reward signals can yield nontrivial gains, particularly when paired with effective exploration strategies.
@@
To improve credit assignment and reduce wasted exploration, several works extend the outcome-only paradigm with denser, step-level signals. DeepSeek-Prover-v2 (DeepSeek-AI et al., 2024) designs a dual-model pipeline to unify both informal (natural-language) and formal (Lean4) mathematical reasoning to reinforce the formal reasoning ability. It introduces subgoal decomposition, where a prover model solves recursively decomposed subgoals and receives binary Lean feedback at the subgoal level, effectively providing denser supervision and improving both accuracy and interpretability. Following this dual-role collaborative mindset, ProofNet++ (Ambati, 2025) implements a neuro-symbolic RL framework featuring a Symbolic Reasoning Interface, which maps LLM-generated reasoning into formal proof trees, and a Formal Verification Engine, which verifies these proofs with Lean or HOL Light and routes error feedback back to the LLM for self-correction. Leanabell-Prover-v2 (Ji et al., 2025) integrates verifier messages into reinforcement updates within a long CoT framework, enabling explicit verifier-aware self-monitoring that stabilizes tactic generation and reduces repeated failure patterns.
@@
Although both outcome-only and process-aware reward paradigms have demonstrated encouraging advances, the scarcity of high-quality theorem-proving data further amplifies the challenges of reinforcement learning under sparse rewards as well as the design of step-level preference signals (Zeng & Zhong, 2024; Wang et al., 2024f; Dong & Ma, 2025). To mitigate these limitations, a prominent line of work adopts expert iteration (ExIt) (Anthony et al., 2017), a framework that combines search with policy learning. This paradigm provides an alternative to outcome-only or process-aware RL, alleviating data scarcity by producing high-quality supervised trajectories. Instead of directly optimizing against sparse verifier signals, ExIt performs search-guided data augmentation: valid proof trajectories discovered by search and checked by a verifier are reused as expert demonstrations in an imitation-learning loop. It usually employs a two-role system: the expert collects valid and progressive trajectories via MCTS under outcome-only verifier feedback, while the apprentice trains a policy on these process-level trajectories and then shares the improved policy back with the expert, thereby bootstrapping subsequent rounds of search and accelerating convergence. Prior work (Polu & Sutskever, 2020) introduces ExIt into formal theorem proving, demonstrating that search-generated expert data can bootstrap models toward tackling complex multi-step proving challenges. Later works adapt this design to Lean and other ITPs.
@@
When applied to formal theorem proving, naive tree search methods often face severe search space explosion when navigating the vast parameterized tactic space. To mitigate this, InternLM2.5-StepProver (Wu et al., 2025m) introduces a preference-based critic model, trained with RLHF-style optimization, to guide expert search, effectively providing a curriculum that directs exploration toward problems of suitable difficulty. Lean-STaR (Lin et al., 2025a) further enhances ExIt by integrating Self-Taught Reasoner (STaR) (Zelikman et al., 2022). It first trains a thought-augmented tactic predictor on synthesized (proof state, generated thought, ground-truth tactic) triples. Then, in the expert-iteration loop, the model produces trajectories that interleave thoughts with tactics; trajectories with tactics successfully validated by Lean are retained and reused for imitation learning. Empirically, the inclusion of thoughts increases the diversity of exploration in the sample-based proof search.
@@
A recent work, STP (Dong & Ma, 2025), points out that solely relying on expert iteration will quickly plateau due to the sparse positive rewards. To address this, it extends the conjecturer–prover self-play idea from Minimo (Poesia et al., 2024) to practical formal languages (Lean/Isabelle) with an open-ended action space and starts from a pretrained model. STP instantiates a dual-role loop in which a conjecturer proposes statements that are barely provable by the current prover, and a prover is trained with standard expert iteration; this generates an adaptive curriculum and alleviates sparse training signals. Empirically, STP reports large gains on LeanWorkbook (Ying et al., 2025) and reports competitive results among whole-proof generation methods on miniF2F (Zheng et al., 2022) and ProofNet (Azerbayev et al., 2023).
Figure 8: Benchmark Performance of RL-Enhanced Math Methods. Scores are pass@1 unless otherwise specified.
@@
ARTIST (Singh et al., 2025) Outcome -
ToRL (Li et al., 2025l) Outcome GitHub [Uncaptioned image]HuggingFace
ZeroTIR (Mai et al., 2025) Outcome GitHub [Uncaptioned image]HuggingFace
TTRL (Zuo et al., 2025) Outcome GitHub
RENT (Prabhudesai et al., 2025) Outcome GitHub Website
Satori (Shen et al., 2025b) Outcome GitHub [Uncaptioned image]HuggingFace Website
1-shot RLVR (Wang et al., 2025r) Outcome GitHub [Uncaptioned image]HuggingFace
Prover-Verifier Games  (Kirchner et al., 2024) Outcome -
rStar2-Agent (Shang et al., 2025a) Outcome GitHub
START (Li et al., 2025b) Process -
LADDER (Simonds & Yoshiyama, 2025) Process -
SWiRL (Goldie et al., 2025) Process -
RLoT (Hao et al., 2025b) Process GitHub
@@
DeepSeek-Prover-v1.5 (Xin et al., 2025) Outcome GitHub [Uncaptioned image]HuggingFace
Leanabell-Prover (Zhang et al., 2025k) Outcome GitHub [Uncaptioned image]HuggingFace
Kimina-Prover  (Wang et al., 2025a) Outcome GitHub [Uncaptioned image]HuggingFace
Seed-Prover (Chen et al., 2025c) Outcome GitHub
DeepSeek-Prover-v2 (DeepSeek-AI et al., 2024) Process GitHub [Uncaptioned image]HuggingFace
ProofNet++ (Ambati, 2025) Process -
Leanabell-Prover-v2 (Ji et al., 2025) Process GitHub
InternLM2.5-StepProver (Wu et al., 2025m) Hybrid GitHub
Lean-STaR (Lin et al., 2025a) Hybrid GitHub [Uncaptioned image]HuggingFace Website
STP (Dong & Ma, 2025) Hybrid GitHub [Uncaptioned image]HuggingFace
@@ 4.4 GUI Agent
GUI agents have progressed through distinct training paradigms. Early systems used pre-trained vision–language models (VLMs) in a pure zero-shot fashion, mapping screenshots and prompts directly to single-step actions. Later, SFT on static (screen, action) trajectories improved grounding and reasoning, but were limited by scarce human operation traces. Reinforcement fine-tuning (RFT) reframes GUI interaction as sequential decision-making, allowing agents to learn via trial-and-error with sparse or shaped rewards, and has advanced from simple single-task settings to complex, real-world, long-horizon scenarios. Table 7 presents the majority of works studied in this section.
@@ 4.4.1 RL-free Methods
Early GUI agents directly leveraged pre-trained Vision–Language Models (VLMs) in a purely zero-shot manner, mapping screenshots and prompts to single-step actions without any task-specific fine-tuning. Representative systems include MM-Navigator (Yan et al., 2023), SeeAct (Zheng et al., 2024), and TRISHUL (Kunal Singh, 2025), which differ in interface domains or parsing strategies but share the same reliance on off-the-shelf VLMs. While showcasing the generality of foundation models, these approaches suffer from limited grounding accuracy and reliability, restricting their applicability to complex tasks (Zhang et al., 2025b; Nguyen et al., 2025a).
@@
The SFT paradigm adapts pre-trained vision–language models to GUI tasks by minimizing cross-entropy loss on offline (screen, action) pairs, without online interaction. InfiGUIAgent (Liu et al., 2025n) employs a two-stage pipeline that first improves grounding and then incorporates hierarchical and reflective reasoning. UI-AGILE (Lian et al., 2025) enhances supervised fine-tuning by incorporating continuous rewards, simplified reasoning, and cropping-based resampling, while further proposing a decomposed grounding mechanism for handling high-resolution displays. TongUI (Zhang et al., 2025a) instead emphasizes data scale, constructing the 143K-trajectory GUI-Net from multimodal web tutorials to enhance generalization. While differing in focus, these approaches all face the limitation of scarce human operation traces.
@@ 4.4.2 RL in Static GUI Environments
In static settings, reinforcement learning is applied on pre-collected datasets with deterministic execution traces, using rule-based criteria for outcome evaluation in the absence of live environment interactions. GUI-R1 (Luo et al., 2025d) adopts an R1-style reinforcement fine-tuning pipeline over a unified action schema, using simple format and correctness rewards to improve step-level action prediction with modest data. UI-R1 (Lu et al., 2025d) applies group-relative policy optimization to stabilize policy updates and improve exact parameter matching through a compact action interface and reward shaping for action-type and argument accuracy. InFiGUI-R1 (Liu et al., 2025o) introduces a two-stage training paradigm that first distills spatial reasoning to enhance grounding, followed by reinforcement learning with sub-goal supervision and recovery mechanisms to improve long-horizon reasoning. AgentCPM-GUI (Zhang et al., 2025y) combines grounding-aware pre-training, supervised imitation, and GRPO-based reinforcement fine-tuning with a concise JSON action space, reducing decoding overhead while improving robustness on long-horizon sequences. UI-Venus (Gu et al., 2025) is a multimodal screenshot-based UI agent fine-tuned via RFT with custom reward functions and a self-evolving trajectory framework, achieving a new state-of-the-art performance in both UI grounding and navigation.
@@ 4.4.3 RL in Interactive GUI Environments
In interactive settings, reinforcement learning agents are optimized through online rollouts in dynamic environments, requiring robustness to stochastic transitions and long-horizon dependencies. WebAgent-R1 (Wei et al., 2025f) conducts end-to-end multi-turn reinforcement learning with asynchronous trajectory generation and group-wise advantages, improving success on diverse web tasks. Vattikonda et al. (2025) studies reinforcement learning for web agents under realistic page dynamics and large action spaces, highlighting challenges in credit assignment and safe exploration. UI-TARS (Qin et al., 2025) integrates pre-training for GUI understanding with reinforcement learning for native desktop control, coupling milestone tracking and reflection to enhance long-horizon execution. DiGiRL (Bai et al., 2024) introduces an offline-to-online reinforcement learning pipeline on real Android devices, combining advantage-weighted updates, doubly robust advantage estimation, and instruction-level curricula to cope with non-stationarity. ZeroGUI (Yang et al., 2025b) automates task generation and reward estimation with a vision-language evaluator, then applies two-stage online reinforcement learning (training on generated tasks followed by test-time adaptation) to reduce human supervision. MobileGUI-RL (Shi et al., 2025c) scales training on Android virtual devices with trajectory-aware GRPO, a decaying efficiency reward, and curriculum filtering, improving execution efficiency and generalization while keeping the system practical for large rollout volumes. ComputerRL (Lai et al., 2025) introduces an API‑GUI hybrid interaction paradigm paired with a massively parallel, fully asynchronous RL infrastructure and the novel Entropulse training strategy—alternating RL with supervised fine‑tuning—to empower GUI‑based agents to operate efficiently and scalably in desktop environments.
@@
MM-Navigator (Yan et al., 2023) Vanilla VLM - GitHub
SeeAct (Zheng et al., 2024) Vanilla VLM - GitHub
TRISHUL (Kunal Singh, 2025) Vanilla VLM - -
InfiGUIAgent (Liu et al., 2025n) SFT Static GitHub [Uncaptioned image]HuggingFace Website
UI-AGILE (Lian et al., 2025) SFT Interactive GitHub [Uncaptioned image]HuggingFace
TongUI (Zhang et al., 2025a) SFT Static GitHub [Uncaptioned image]HuggingFace Website
@@
GUI-R1 (Luo et al., 2025d) RL Static GitHub [Uncaptioned image]HuggingFace
UI-R1 (Lu et al., 2025d) RL Static GitHub [Uncaptioned image]HuggingFace
InFiGUI-R1 (Liu et al., 2025o) RL Static GitHub [Uncaptioned image]HuggingFace
AgentCPM (Zhang et al., 2025y) RL Static GitHub [Uncaptioned image]HuggingFace
UI-Venus (Gu et al., 2025) RL Static GitHub
WebAgent-R1 (Wei et al., 2025f) RL Interactive -
Vattikonda et al. (2025) RL Interactive -
UI-TARS (Qin et al., 2025) RL Interactive GitHub [Uncaptioned image]HuggingFace Website
UI-TARS-2 (Wang et al., 2025c) RL Interactive GitHub Website
DiGiRL (Bai et al., 2024) RL Interactive GitHub [Uncaptioned image]HuggingFace Website
ZeroGUI (Yang et al., 2025b) RL Interactive GitHub
MobileGUI-RL (Shi et al., 2025c) RL Interactive -
ComputerRL (Lai et al., 2025) RL Interactive -
4.5 Vision Agents
RL has been applied to a wide range of vision tasks (including, but not limited to, image, video, 3D perception and generation). Since the number of related papers is substantial, this section does not aim to provide an exhaustive overview; for a more comprehensive survey on RL for various vision tasks, we refer readers to two dedicated surveys in vision (Wu et al., 2025h; Zhou et al., 2025a).